Circular No. 5714

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY PULSAR IN N63A K. Dennerl and M. Kürster, Max-Planck-Institut für Extraterres-

trische Physik, Garching, communicate: “We report the discovery of pulsed x-ray emission with a barycentric period of 201.0 s from the direction of the supernova remnant (SNR) N63A in the Large Magellanic Cloud (cf. Henize 1956, Ap.J. Suppl. 2, 315) with the ROSAT satellite during a 45 000-s Position Sensitive Proportional Counter (PSPC) observation. Since N63A was located at an off-axis angle of 31′, the spacecraft wobble can be excluded as a source of artificial variability. The pulse profile shows a highly significant modulation consisting of two broad peaks of similar intensity and shape separated by ∼ 0.45 in phase. Due to the degraded Point Spread Function, the observed flux contains contributions from the diffuse x-ray emission of the SNR. The total count rate for the unresolved source region (SNR and pulsar) is modulated between 7.2 and 7.8 cts/s in the band 0.1-2.4 keV, corrected for vignetting. The peak-to-valley ratio of the modulation is increased from 1.1 to 1.5 by a spectral selection of the band 0.7-1.1 keV and a rough spatial selection of the pulsar region based on the pulsed image. The pulsed emission can be described by a blackbody spectrum with a temperature of 0.15 keV. This implies a luminosity of the pulsed flux of 5 × 10^{35} erg/s in the total PSPC energy range if the pulsar is located in the LMC. We have also found this pulsar in the ROSAT all-sky survey data. Further analysis is in progress.”

PERIODIC COMET SCHWASSMANN-WACHMANN 1 The observations by A. Nakamura reported on IAUC 5704 were made

with a CCD. Further total magnitude estimates (cf. IAUC 5708): Feb. 10.14 UT, 12.7 (A. Hale, Las Cruces, NM, 0.41-m reflector; visual); 12.53, 12.5 (T. Kojima, Chiyoda, Gunma-ken, Japan, 0.25-m f/3.4 Wright-Schmidt camera + TP 6415 film); 14.53, 12.6 (A. Nakamura, Kuma, Japan, 0.20-m reflector; visual); 15.93, 12.7 (J.-C. Merlin, Le Creusot, France, 0.40-m reflector; T-Max 400 film + image intensifier); 16.74, 12.8 (H. Mikuž, Ljubljana, Slovenia, 0.20-m f/2 Baker-Schmidt camera + V filter + CCD); 18.29, 12.8 (Hale); 20.21, 12.5 (I. Ferrin and F. Fuenmayor, Mérida, Venezuela, 1-m Schmidt telescope + red-sensitive CCD); 20.78, 13.4 (Mikuž).

1993 February 25 Daniel W. E. Green